OpenStudy (anonymous):

Help please? Solve: log2 (x+5) + log2 (x+1)=5

OpenStudy (anonymous):

@Hero can you help me out

hero (hero):

\[log_a(x) + \log_a(y) = \log_a(xy)\]

OpenStudy (anonymous):

I got an answer 3 but I want to know if I am correct

OpenStudy (anonymous):

@hero

hero (hero):

Show your work

OpenStudy (anonymous):

@Hero log2(x+5)(x+1)=5 log2(x^2 +6x+5)=5 2^5=x^2+6x+5 32=x^2+6x+5 0=x^2+6x-27 0=(x-3)(x+9) x=3 & -9 solution is 3 because you can't have a -9

OpenStudy (anonymous):

I was talking to someone and they said there was no solution that is why I was worried

hero (hero):

x = 3 is the correct solution

OpenStudy (anonymous):

so there is a solution thank God I did it correct then thank you so much for confirming

hero (hero):

Make sure to use double parentheses for the first step: log_{2}((x+5)(x+1))=5